A population consist of the four numbers 1, 2, 4 and 5. List all the possible samples of size n=3 which can be drawn with replacement from population


1
Expert's answer
2021-06-28T16:42:06-0400

We have population values "1,2,4,5" population size "N=4" and sample size "n=3."Thus, the number of possible samples which can be drawn with replacement is


"(N)^n=4^3=64"


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample & Sample \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,1,1 \\
 \hdashline
 2 & 1,1,2 \\
 \hdashline
 3 & 1,1,4 \\
 \hdashline
 4 & 1,1.5 \\
 \hdashline
 5 & 1,2,1 \\
\hdashline
 6 & 1,2,2 \\
 \hline
 7 & 1,2,4 \\
 \hline
 8 & 1,2,5 \\
 \hline
 9 & 1,4,1 \\
 \hline
 10 & 1,4,2 \\
 \hline
 11 & 1,4,4 \\
 \hline
12 & 1,4,5 \\
 \hline
13 & 1,5,1 \\
 \hline
14 & 1,5,2 \\
 \hline
15 & 1,5,4 \\
 \hline
16 & 1,5,5 \\
 \hline
17 & 2,1,1 \\
 \hdashline
 18 & 2,1,2 \\
 \hdashline
 19 & 2,1,4 \\
 \hdashline
 20 & 2,1.5 \\
 \hdashline
 21 & 2,2,1 \\
\hdashline
 22 & 2,2,2 \\
 \hline
 23 & 2,2,4 \\
 \hline
 24 & 2,2,5 \\
 \hline
 25 & 2,4,1 \\
 \hline
 26 & 2,4,2 \\
 \hline
 27 & 2,4,4 \\
 \hline
28 & 2,4,5 \\
 \hline
29 & 2,5,1 \\
 \hline
30 & 2,5,2 \\
 \hline
31 & 2,5,4 \\
 \hline
32 & 2,5,5 \\
 \hline
33 & 4,1,1 \\
 \hdashline
 34 & 4,1,2 \\
 \hdashline
 35 & 4,1,4 \\
 \hdashline
 36 & 4,1.5 \\
 \hdashline
 37 & 4,2,1 \\
\hdashline
 38 & 4,2,2 \\
 \hline
 39 & 4,2,4 \\
 \hline
 40 & 4,2,5 \\
 \hline
 41 & 4,4,1 \\
 \hline
 42 & 4,4,2 \\
 \hline
 43 & 4,4,4 \\
 \hline
44 & 4,4,5 \\
 \hline
45 & 4,5,1 \\
 \hline
46 & 4,5,2 \\
 \hline
47 & 4,5,4 \\
 \hline
48 & 4,5,5 \\
 \hline
49 & 5,1,1 \\
 \hdashline
 50 & 5,1,2 \\
 \hdashline
 51 & 5,1,4 \\
 \hdashline
 52 & 5,1.5 \\
 \hdashline
 53 & 5,2,1 \\
\hdashline
 54 & 5,2,2 \\
 \hline
 55 & 5,2,4 \\
 \hline
 56 & 5,2,5 \\
 \hline
 57 & 5,4,1 \\
 \hline
 58 & 5,4,2 \\
 \hline
 59 & 5,4,4 \\
 \hline
60 & 5,4,5 \\
 \hline
61 & 5,5,1 \\
 \hline
62 & 5,5,2 \\
 \hline
63 & 5,5,4 \\
 \hline
64 & 5,5,5 \\
 \hline
\end{array}"